# DeferredGS: Decoupled and Editable Gaussian Splatting with Deferred Shading

### v1

@@ 1 Introduction
Reconstructing and editing 3D scenes from multi-view images is an important task in computer graphics and computer vision. Recently, Neural radiance fields (NeRFs) [1] have shown promising results in novel view synthesis, based on which a number of editing methods [2, 3, 4, 5] are proposed. However, NeRF-based methods require dense sampling of camera rays, making visualization less efficient. Gaussian splatting [6] instead approximates a scene with a set of 3D Gaussians and visualizes it with a rasterization-based renderer to achieve real-time performance. However, despite high-quality and efficient rendering, Gaussian splatting still lacks the ability of editing.
In the original Gaussian splatting representation, the appearance is approximated by a view-related spherical harmonic function which entangles both texture and lighting information, making separate texture and lighting editing impossible. Independent editing requires texture and lighting decomposition so that one part can be manipulated without influencing the other part. However, different from 3D representations like textured meshes or NeRFs that have an explicit surface or a continuous implicit field, which can provide a good geometry basis like normal direction for the texture and lighting decomposition, Gaussian splatting models a scene’s geometry with a set of discrete Gaussians, and this discrete representation makes the geometry less tractable, posing challenges for later texture and lighting decoupling.
@@
Recently, in the context of Gaussian splatting, a few methods have made attempts to edit it. Two independent works both named GaussianEditor [7, 8] utilize the 2D segmentation model [9] and text-to-image diffusion models [10] to enable scene editing like object removal/insertion or text-guided modification similar to Instruct-NeRF2NeRF [11]. To enable more tractable editing on texture and lighting, the works [12, 13, 14, 15] try to disentangle texture and lighting to enable individual editing for both diffuse scenes and reflective scenes. Although obtaining plausible decomposition results, they use forward shading that first calculates shading for each Gaussian and blends shaded Gaussians, which can cause blending artifacts when relighting a scene. This is because the geometry attributes of Gaussians like opacity are optimized under the input lighting condition while under novel illuminations shaded Gaussians can be blended in a wrong way (see Fig. 5 for some examples). In addition, they learn geometry from the Gaussian splatting representation with the normal direction derived from the rendered depth map so their geometry may contain noise and degenerate on scenes with reflective surfaces (see Fig. 4).
To resolve the aforementioned issues, we propose DeferredGS that decouples a scene’s texture and its surrounding lighting of the Gaussian splatting representation with the deferred shading technique. Given multi-view captures of an input scene, we define a set of Gaussians with extra learnable texture attributes and normal directions attached to model texture and geometry. To enable faithful geometry reconstruction, we distill the normal field from a signed distance function (SDF) onto Gaussians. We also optimize a learnable environment map to approximate the surrounding lighting conditions. In the rendering process, geometry and texture buffers are first rasterized and shading is calculated at pixel level with the deferred shading technique to enable faithful decoupling and editing. Though involving more complex shading computation, DeferredGS still allows real-time rendering ($\sim$30FPS at $800\times 800$ resolution on a single 3090 GPU).
@@
  • We propose DeferredGS, a decoupled Gaussian splatting representation with editable geometry, texture and lighting, where its geometry is enhanced by a normal distillation module.
@@ 2.1 Neural Rendering
With the development of recent implicit representations [16, 17, 18] and volume rendering [19], neural rendering has become popular in 3D scene reconstruction. The pioneering work Neural Radiance Field (NeRF) [1] models a sample point’s density and view-dependent color with two neural networks and renders the scene by integrating sample points’ colors along the rays. NeRF can synthesize realistic novel view results but dense sampling and network query make both training and rendering slow. To improve the efficiency, a few hybrid rendering methods bake view-independent features onto regular or irregular geometry proxies. NSVF [20] models 3D scenes with an octree structure and prunes voxels with low-density values so that sample points outside the octree can be ignored in the rendering process. DVGO [21] stores density values and appearance features on two voxel grids and directly optimizes these two voxel grids to reconstruct 3D scenes. TensorRF [22] models a scene with a set of matrices and vector pairs that depict the scene’s geometry and appearance on corresponding planes, which not only accelerates the training speed but also reduces the storage size. Instant-NGP [23] instead utilizes multi-resolution hash encoding as its scene representation and enables training convergence in as little as 5 seconds. MobileNeRF [24] and BakedSDF [25] share a similar idea to first reconstruct the mesh of a scene and attach learnable appearance information onto mesh vertices, which enables real-time rendering on common devices. Moving a step forward, Gaussian splatting [6] abandons the neural network and models a 3D scene as a set of 3D Gaussians on which attributes like position, rotation, scaling, opacity, and spherical harmonic coefficients are defined. Gaussian splatting uses a rasterization-based renderer which can produce rendered results from a viewpoint in real time. Despite the high-quality novel view results, these methods still lack the capability of editing, making appearance change or relighting impossible.
@@ 2.3 Inverse Rendering
The goal of inverse rendering is to recover geometry, material and lighting information from a set of images of a single scene under the same or different illuminations. NeRV [37] is the pioneering work that decomposes geometry and BRDF material under a given lighting condition. To reduce dependence on light input, later methods work under unknown light conditions. NeRD [38] and PhySG [39] use Spherical Gaussian (SG) as the light representation and utilize a density field and a signed distance field as their geometry representations. Based on PhySG, InvRender [40] further takes visibility into consideration to better recover material information in occluded regions. NeROIC [41] approximates the illumination with Spherical Harmonic functions and enables training on Internet images under different lighting conditions. NeRFactor [42] represents the lighting with a low-resolution image. To better estimate material from images, it also trains a BRDF autoencoder. NDR [43] and NDRMC [44] use a hybrid geometry representation DMTet [45] to enable efficient inverse rendering. NMF [46] and TensoIR [47] introduce the importance sampling strategy for effective shading calculation. To model reflective objects, NeRO [48] and DE-NeRF [5] express high-frequency lighting as MLP networks. Apart from works concentrating on object-level scenes, there are methods [49, 50, 51] targeting large-scale outdoor scenes. There are also concurrent inverse rendering works based on the Gaussian splatting representation like [12, 13, 15, 14]. Different from them, our method introduces a normal distillation module to enable more accurate normal estimation and alleviate blending artifacts with the deferred shading technique when relighting Gaussians.
@@ 3 Method
We propose DeferredGS, a decoupled Gaussian splatting representation that decomposes geometry, texture, and lighting from multi-view images, which enables downstream applications like scene relighting and texture editing. The overview of our method is shown in Fig. 2. Our ultimate goal is to extract geometry, texture, and lighting information from a set of captured images. First, we briefly introduce the rendering formulation in Gaussian splatting and define the texture parameters we use in rendering (Sec. 3.1). Since directly optimizing everything from scratch can lead to suboptimal results, especially on the geometry side, we introduce a normal field distillation module (Sec. 3.2) that utilizes an SDF network to guide Gaussians to a better surface representation. With geometry set up properly, we bind texture parameters onto the Gaussian splatting representation and perform deferred shading to realize plausible decomposition (Sec. 3.3). Finally, this decomposition enables scene editing applications like relighting and texture editing (Sec. 3.4).
@@ 3.1 Gaussian Splatting and Texture Definition
Gaussian splatting models a 3D scene by a set of Gaussians with which attributes like position $P$, rotation $R$, scale $S$, opacity $\alpha$ and Spherical Harmonic coefficients (SH) representing view-dependent appearance are associated. It uses a point-based $\alpha$-blending rendering formulation similar to NeRF [1].
@@
where $N$ is the number of sample points on a ray. $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is accumulated transmittance and * can be any attribute of the sample point to be blended such as depth, color and normal direction. Instead of densely sampling on the ray, Gaussian splatting efficiently projects 3D Gaussians onto a 2D image plane and approximates 3D scenes by blending view-dependent color represented by SH. Instead of modeling the appearance as a whole with SH, our DeferredGS replaces SH with optimizable texture parameters including diffuse albedo $k_{d}$, roughness $r$, and specular albedo $k_{s}$ as shown in Fig. 2 to enable texture and lighting decoupling.
@@ 3.2 Normal Field Distillation
where $V$ is the number of rays in a batch. $M$ is the number of sample points on a single ray and $C^{t}(v)$ is the corresponding ground truth color. The second term is the Eikonal loss, where $\|\nabla_{x_{v,i}}\|$ is the spatial norm of the SDF network $f_{s}(x)$’s gradient at the $i$th sample point $x_{v,i}$ on the ray $v$. In order to distill the normal information from Instant-RefNeuS to Gaussian splatting, we define an extra normal direction $n$ on each Gaussian as shown in Fig. 2 and render the normal of Gaussian splatting for a pixel $C_{n}$ using Eqn. (1). Then we distill the normal field Gaussian splatting by minimizing the difference between $C_{n}$ and the corresponding rendered normal in Instant-RefNeuS:
@@ 3.3 Deferred Shading in Gaussian Splatting
where $L(\omega_{o})$ and $L_{i}(\omega_{i})$ represent outgoing lighting in direction $\omega_{o}$ and incident lighting from direction $\omega_{i}$ respectively. $f(\omega_{i},\omega_{o})$ is the BRDF properties at a point which can be determined by the Disney shading model [53]:
@@
We model the scene’s environment lighting with a $6\times 512\times 512$ High Dynamic Range (HDR) cube map and utilize the SplitSum [55] approximation which separates the integrals of lighting and BRDF to enable efficient shading calculation. For the diffuse color $c_{dif}$:
@@
where $L_{i}(\omega_{i})(\omega_{i}\cdot n)$ is only dependent on the normal direction $n$ and can be pre-computed and stored in a 2D texture.
@@
where $Int_{light}$ represents the integral of incident light with the normal distribution function $D$. Given a fixed environment map, $Int_{light}$ is only related to the roughness value $r$ and therefore can be pre-computed and stored in a mipmap where each mip level corresponds to a fixed roughness value. $Int_{BRDF}$ is the integral of specular BRDF under a uniform white environment lighting. It is determined by the roughness value in the BRDF and the dot product between incident light direction and normal direction $\omega_{i}\cdot n$. Again it can be pre-computed and stored in a 2D texture and efficiently looked up in rendering. The final shaded color is $c_{g}=c_{dif}+c_{spec}$.
A straightforward way to obtain a decoupled Gaussian splatting representation is performing forward shading for each Gaussian to get its color $c_{i}$ and rasterizing Gaussians with Eqn. (1) similar to previous works [40, 42, 15, 14]. Though obtaining plausible decoupled results, we observe notable artifacts when rendering the decoupled Gaussians under a novel lighting condition (see Fig. 8 and Fig. 5). This is because even though Eqn. (3) constrains the rasterized normal to be close to that of Instant-RefNeuS at pixel level, the rasterized normal is still blended by multiple Gaussians’ normals. The geometry attributes like position, covariance and opacity may overfit the given lighting condition to produce proper blended normal and shaded color and it does not guarantee that each Gaussian’s individual normal or shaded color is correct. Thus when changing the lighting condition, the trained geometry attributes may not work for the shaded colors under a new lighting condition and blend them in a wrong way.
@@ 4.1 Datasets and Metrics
We conduct qualitative and quantitative experiments on two synthetic datasets, NeRF Synthetic [1] and Shiny Blender [34] and one real dataset, the Stanford ORB dataset  [58]. For all datasets, we evaluate the DeferredGS and baseline methods on novel view synthesis, decomposition and relighting tasks. For quantitative comparisons, we use PSNR, SSIM [59] and LPIPS [60] metrics to compare the similarity between images. To evaluate the quality of the estimated normals, we compare estimated normal images with the ground truth with Mean Squared Error (MSE).
@@ 4.2 Implementation Details
We train the decoupled Gaussian splatting representation on a single 3090 GPU with 24GB memory. The training process takes 3-4 hours depending on the scenes’ complexity and the input images’ resolution. DeferredGS allows $\sim$30FPS novel view synthesis at $800\times 800$ resolution on a single 3090 GPU. For hyperparameter settings, in Eqn. (2) of the main paper, $\lambda_{1}$ is set to $0.1$. In Eqn. (8) of the main paper, $\lambda_{nd},\lambda_{L1},\lambda_{ssim},\lambda_{mask},\lambda_{TV}$ are set to $0.5,0.8,0.2,0.5,0.001$ in all our examples.
@@ 4.3 Novel View Synthesis
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption

### v2

@@ 1 Introduction
Reconstructing and editing 3D scenes from multi-view images is an important task in computer graphics and computer vision. Recently, Neural Radiance Fields (NeRFs) [1] have shown promising results in novel view synthesis, based on which a number of editing methods [2, 3, 4, 5] are proposed. However, NeRF-based methods require dense sampling of camera rays, making rendering less efficient. Gaussian splatting [6] instead approximates a scene with a set of 3D Gaussians and visualizes it with a rasterization-based renderer to achieve real-time performance. However, despite high-quality and efficient rendering, Gaussian splatting still lacks the ability of editing.
In the original Gaussian splatting representation, the appearance is approximated by a view-related spherical harmonic function which entangles both texture and lighting information, making separate texture and lighting editing impossible. Independent editing requires texture and lighting decomposition so that one part can be manipulated without influencing the other. However, different from 3D representations like textured meshes or NeRFs that have an explicit surface or a continuous implicit field, which can provide a good geometry basis like normal direction for the texture and lighting decomposition, Gaussian splatting models a scene’s geometry with a discrete set of Gaussians, and this discrete representation makes the geometry less tractable, posing challenges for later texture and lighting decoupling.
@@
Recently, in the context of Gaussian splatting, a few methods have made attempts to edit it. Two independent works both named GaussianEditor [7, 8] utilize the 2D segmentation model [9] and text-to-image diffusion models [10] to enable scene editing like object removal/insertion or text-guided modification similar to Instruct-NeRF2NeRF [11]. To enable better-controlled editing on texture and lighting, the works [12, 13, 14, 15] try to disentangle texture and lighting to enable individual editing for both diffuse scenes and reflective scenes. Although obtaining plausible decomposition results, they use forward shading that first calculates shading for each Gaussian and then blends shaded Gaussians, which can cause blending artifacts when the scene is relighted. This is because the geometry attributes of Gaussians like opacity are optimized under the input lighting condition while under novel illuminations shaded Gaussians might be blended in a wrong way (see 3rd to 6th rows in Fig. 5 for some examples). In addition, they learn geometry from the Gaussian splatting representation with the normal direction derived from the rendered depth map so their geometry may contain noise and degenerate on scenes with reflective surfaces (see the “Normal” rows in Fig. 4).
To resolve the aforementioned issues, we propose DeferredGS that decouples a scene’s texture and its surrounding illumination of the Gaussian splatting representation with the deferred shading technique. Given multi-view captures of an input scene, we define a set of Gaussians with extra learnable texture attributes and normal directions attached to model texture and geometry respectively. To enable faithful geometry reconstruction, we distill the normal field from a signed distance function (SDF) onto the Gaussians. We also optimize a learnable environment map to approximate the surrounding lighting conditions. In the rendering process, geometry and texture buffers are first rasterized and shading is calculated at the pixel level with the deferred shading technique to enable faithful decoupling and editing. Despite involving more complex shading computation, DeferredGS still allows real-time rendering ($\sim$30FPS at $800\times 800$ resolution on a single 3090 GPU).
@@
  • We introduce DeferredGS, a decoupled Gaussian splatting representation with editable geometry, texture and lighting, where its geometry is enhanced by a normal distillation module.
@@ 2.1 Neural Rendering
With the development of recent implicit representations [16, 17, 18] and volume rendering [19], neural rendering has become popular in 3D scene reconstruction. The pioneering work Neural Radiance Field (NeRF) [1] models a sample point’s density and view-dependent color with two neural networks and renders the scene by integrating sample points’ colors along the rays. NeRF can synthesize realistic novel view results but dense sampling and network query make both training and rendering slow. To improve the efficiency, a few hybrid rendering methods bake view-independent features onto regular or irregular geometry proxies. NSVF [20] models 3D scenes with an octree structure and prunes voxels with low-density values so that sample points outside the octree can be ignored in the rendering process. DVGO [21] stores density values and appearance features on two voxel grids and directly optimizes these two voxel grids to reconstruct 3D scenes. TensoRF [22] models a scene with a set of matrices and vector pairs that depict the scene’s geometry and appearance on corresponding planes, which not only accelerates the training speed but also reduces the storage size. Instant-NGP [23] instead utilizes multi-resolution hash encoding as its scene representation and enables training convergence in as little as 5 seconds. MobileNeRF [24] and BakedSDF [25] share a similar idea to first reconstruct the mesh of a scene and attach learnable appearance information onto mesh vertices, which enables real-time rendering on consumer devices. Moving a step forward, Gaussian splatting [6] abandons the neural network and models a 3D scene as a set of 3D Gaussians on which attributes like position, rotation, scaling, opacity, and spherical harmonic coefficients can be tuned. Gaussian splatting uses a rasterization-based renderer which can produce rendered results from a viewpoint in real time. To better reconstruct reflective scenes, GaussianShader [14] proposes to model view-dependent effects separately similar to RefNeRF [26] and Ye et al. [27] further introduce deferred rendering for reflection modeling. Despite the high-quality novel view results, these methods still lack the capability of editing, making appearance change or relighting impossible.
@@ 2.3 Inverse Rendering
The goal of inverse rendering is to recover geometry, material and lighting information from a set of images of a single scene under the same or different illuminations. NeRV [38] is the pioneering work that decomposes geometry and BRDF (Bidirectional Reflectance Distribution) material under a given lighting condition. To reduce dependence on light input, later methods work under unknown light conditions. NeRD [39] and PhySG [40] use Spherical Gaussian (SG) as the light representation and utilize a density field and a signed distance field as their geometry representations. Based on PhySG, InvRender [41] further takes visibility into consideration to better recover material information in occluded regions. NeROIC [42] approximates the illumination with Spherical Harmonic functions and enables training on Internet images under different lighting conditions. NeRFactor [43] represents the lighting with a low-resolution image. To better estimate material from images, it also trains a BRDF autoencoder. NDR [44] and NDRMC [45] use a hybrid geometry representation DMTet [46] to enable efficient inverse rendering. NMF [47] and TensoIR [48] introduce the importance sampling strategy for effective shading calculation. To model reflective objects, NeRO [49] and DE-NeRF [5] express high-frequency lighting as MLP networks. Apart from works concentrating on object-level scenes, there are methods [50, 51, 52] targeting large-scale outdoor scenes. There are also concurrent inverse rendering works based on the Gaussian splatting representation like [12, 13, 15, 14]. Different from them, our method introduces a normal distillation module to enable more accurate normal estimation and alleviate blending artifacts with the deferred shading technique when relighting Gaussians.
@@ 3 Method
We propose DeferredGS, a decoupled Gaussian splatting representation that decomposes geometry, texture, and lighting from multi-view images, which enables downstream applications like scene relighting and texture editing. The overview of our method is shown in Fig. 2. First, we briefly introduce the rendering formulation in Gaussian splatting and define the texture parameters we use in rendering (Sec. 3.1). Since directly optimizing everything from scratch can lead to suboptimal results, especially on the geometry side, we introduce a normal field distillation module (Sec. 3.2) that utilizes an SDF network to guide Gaussians to a better surface representation. With geometry set up properly, we bind texture parameters onto the Gaussian splatting representation and perform deferred shading to obtain plausible decomposition (Sec. 3.3). Finally, we demonstrate how this decomposition enables scene editing applications like relighting and texture editing (Sec. 3.4).
@@ 3.1 Gaussian Splatting and Texture Definition
Gaussian splatting models a 3D scene by a set of Gaussians with which attributes like position $P$, covariance $\Sigma$, scale $S$, opacity $\alpha$ and Spherical Harmonic coefficients (SH) representing view-dependent appearance are associated. It uses a point-based $\alpha$-blending rendering formulation similar to NeRF [1].
@@
where $N$ is the number of sample points on a ray. $T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j})$ is accumulated transmittance and * can be any attribute of the sample point to be blended such as depth, color and normal direction. Gaussian splatting avoids densely sampling on the ray and efficiently projects 3D Gaussians onto a 2D image plane and approximates 3D scenes by blending view-dependent color represented by SH. Instead of modeling the appearance as a whole with SH, our DeferredGS replaces SH with optimizable texture parameters including diffuse albedo $k_{d}$, roughness $r$, and specular albedo $k_{s}$ as shown in Fig. 2 to enable texture and lighting decoupling.
@@ 3.2 Normal Field Distillation
where $V$ is the number of rays in a batch. $M$ is the number of sample points on a single ray and $C^{t}(v)$ is the corresponding ground truth color. The second term is the Eikonal loss, where $\|\nabla f_{s}({x_{v,i})}\|$ is the spatial norm of the SDF network $f_{s}(x)$’s gradient at the $i$th sample point $x_{v,i}$ on the ray $v$. In order to distill the normal information from Instant-RefNeuS to Gaussian splatting, we define an extra normal direction $n$ on each Gaussian as shown in Fig. 2 and render the normal $C_{n}$ of Gaussian splatting for a pixel using Eqn. (1). Then we distill the normal field Gaussian splatting by minimizing the difference between $C_{n}$ and the corresponding rendered normal in Instant-RefNeuS:
@@ 3.3 Deferred Shading in Gaussian Splatting
where $L_{o}(\omega_{o})$ and $L_{i}(\omega_{i})$ represent outgoing lighting in direction $\omega_{o}$ and incident lighting from direction $\omega_{i}$ respectively. $f(\omega_{i},\omega_{o})$ is the BRDF at a point which can be parameterized by the Disney shading model [54]:
@@
We model the scene’s environment lighting with a $6\times 512\times 512$ High Dynamic Range (HDR) cube map and utilize the SplitSum [56] approximation which separates the integrals of lighting and BRDF to enable efficient shading calculation. T he diffuse color $c_{diff}$:
@@
which is only dependent on the normal direction $n$ and can be pre-computed and stored in a 2D texture.
@@
where $Int_{light}$ represents the integral of incident light with the normal distribution function $D$. Given a fixed environment map, $Int_{light}$ is only related to the roughness value $r$ and therefore can be pre-computed and stored in a mipmap where each mip level corresponds to a fixed roughness value. $Int_{BRDF}$ is the integral of specular BRDF under a uniform white environment lighting. It is determined by the roughness value in the BRDF and the dot product between incident light direction and normal direction $\omega_{i}\cdot n$. Again it can be pre-computed and stored in a 2D texture and efficiently looked up in rendering. The final shaded color is $c_{g}=c_{diff}+c_{spec}$.
A straightforward way to obtain a decoupled Gaussian splatting representation is performing forward shading for each Gaussian to get its color $c_{i}$ and rasterizing Gaussians with Eqn. (1) similar to previous works [41, 43, 15, 14]. Though obtaining plausible decoupled results, we observe notable artifacts when rendering the decoupled Gaussians under a novel lighting condition (see “F.S.” columns in Fig. 8 and “RGS” and “GShader” columns in Fig. 5). This is because even though Eqn. (3) constrains the rasterized normal to be close to that of Instant-RefNeuS at pixel level, the rasterized normal is still blended from multiple Gaussians’ normals. The geometry attributes like position, covariance and opacity may overfit the given lighting condition to produce proper blended normal and shaded color and it does not guarantee that each Gaussian’s individual normal or shaded color is correct. Thus when changing the lighting condition, the trained geometry attributes may not work for the shaded colors under a new lighting condition and blend them in a wrong way.
@@ 4.1 Datasets and Metrics
We conduct qualitative and quantitative experiments on two synthetic datasets, NeRF Synthetic [1] and Shiny Blender [26] and one real dataset, the Stanford ORB dataset  [59]. For all datasets, we evaluate DeferredGS and baseline methods on novel view synthesis, decomposition and relighting tasks. For quantitative comparisons, we use PSNR (Peak Signal-to-Noise Ratio), SSIM (Structural Similarity) [60] and LPIPS (Learned Perceptual Image Patch Similarity) [61] metrics to compare the similarity between images. To evaluate the quality of the estimated normals, we compare the estimated normal images with the ground truth with Mean Squared Error (MSE).
@@ 4.2 Implementation Details
We train the decoupled Gaussian splatting representation on a single 3090 GPU with 24GB memory. The training process takes 3-4 hours depending on the scenes’ complexity and the input images’ resolution. DeferredGS allows $\sim$30FPS novel view synthesis at $800\times 800$ resolution on a single 3090 GPU. For hyperparameter settings, in Eqn. (2), $\lambda_{1}$ is set to $0.1$. In Eqn. (8), $\lambda_{nd},\lambda_{L1},\lambda_{ssim},\lambda_{mask},\lambda_{TV}$ are set to $0.5,0.8,0.2,0.5,0.001$ in all our examples.
@@ 4.3 Novel View Synthesis
Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to captionRefer to caption Refer to caption Refer to captionRefer to caption Refer to captionRefer to caption